Lreynolds24:

Solve the following equation with fraction coefficients by multiplying by the common denominator. Show your work and give your answer as an improper fraction if needed. 5/3r + 4/3 = -8/9 common denominator answer

Timmyspu:

Ok so first what do you know about how to solve the equation.

supie:

We are solving 5/3r + 4/3 = -8/9 right?

Timmyspu:

Yes

supie:

Ok well \(\frac{5}{3}r+\frac{4}{3}={−8}{9}\) subtract \(\frac{4}{3}\) from both sides \(\frac{5}{3}r+\frac{4}{3}−\frac{4}{3}=\frac{−8}{9}−\frac{4}{3}\) \(\frac{5}{3}r=\frac{−20}{9}\) multiply both sides: \(\frac{3}{5}\) \((\frac{3}{5})(\frac{5}{3}r)=(\frac{3}{5})(\frac{−20}{9})\) \( r=\frac{?}{?}\)

supie:

Does that help @Lreynolds24?

Lreynolds24:

a little thanks @supie

supie:

np